OpenStudy (anonymous):

What is tan theta+sec theta=1?

OpenStudy (anonymous):

are you looking for what theta is?

OpenStudy (anonymous):

no its sys solve the eqution on the interval 0<theta(sign)<2pie

OpenStudy (anonymous):

*says

OpenStudy (anonymous):

yea you need the values of theta that will give you 1

OpenStudy (anonymous):

the choices are 0, 5pie/4,pie/4 and no solution

OpenStudy (tkhunny):

First, "pie" is for eating. "pi" is for math. \(\tan(\theta) + \sec(\theta) = 1\) \(\dfrac{\sin(\theta)}{\cos(\theta)} + \dfrac{1}{\cos(\theta)} = 1\) Notice that \(\cos(\theta) = 0\) is NOT acceptable, no matter what else we do. \(\dfrac{\sin(\theta)+1}{\cos(\theta)} = 1\) \(\sin(\theta) + 1 = \cos(\theta)\) \(\sin(\theta) - \cos(\theta) + 1 = 0\) There are a couple of ways to go from here. What do you suggest? Please show your work.